OpenStudy (anonymous):

User: Tyler Switzer In Course: Algebra I V14 ( 3644) Instructor: Patricia Stefani WARNING: You must not leave this exam form! If you try to click back into this exam again prior to submitting, your access to it will be denied! 10.20 Segment Two Exam Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Instructions: IMPORTANT: Read each problem carefully. Make sure you read all the directions before marking your answers. For Fill in the blank items, you may type in any d

OpenStudy (anonymous):

[3.02] The set W represents multiples of 4 to 40. W = {4, 8, 12, 16, 20, 24, 28, 32, 36, 40} The set C represents multiples of 6 to 48. C = {6, 12, 18, 24, 30, 36, 42, 48} How many elements are in the set W intersection C?

OpenStudy (anonymous):

sorry tht was the wrong thing lol

OpenStudy (ghazi):

just check out the common elements between these two sets....that will be your total number of elements

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

3 of them?

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

[4.06]Which equation represents the line passing through the points (1, 1) and (−2, 13)? 4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (ghazi):

just put the values of x and y ...and check out which of the equation is satisfied

OpenStudy (anonymous):

[2.05]Marlene is a sales associate at a department store. Each week she earns $100 plus a 5% commission of all her sales. This week she earned $160. How much, in dollars, were Marlene's sales this week? (Do not put a $ in your answer.)

OpenStudy (anonymous):

im so sorry there are only 4 more

OpenStudy (ghazi):

it's okay

OpenStudy (anonymous):

u get my msg?

OpenStudy (ghazi):

yea

OpenStudy (anonymous):

u dont have to i have a little but hes gettting all pissy

OpenStudy (anonymous):

7.04] A glass cylinder is completely filled with 32m6n2 cubic inches of salt solution. The jug has a base area of 8m3n2 square inches. The height of the jug, in inches, is represented by the expression The quantity of 32 m to the sixth n squared all over 8 m cubed n squared. Which of the following simplified expressions represents the height of the jug in inches?

OpenStudy (anonymous):

32m^6n^2/8m^3n^2

OpenStudy (swissgirl):

hmmm Can i take this exam too lol

OpenStudy (anonymous):

ummmm i guess if u want i just need help

OpenStudy (anonymous):

looks like a test I don't want to partake in...

OpenStudy (anonymous):

i just need help in tht one question

OpenStudy (anonymous):

it's a test, not a homework.

OpenStudy (anonymous):

come on please its timed

OpenStudy (anonymous):

not really you can sit there all day and leave it open and it wont close the others r timed though

OpenStudy (anonymous):

its the last test and school starts tommarow

OpenStudy (anonymous):

[7.03] A film set designer is using white and colored tiles in a pattern to create paths of different lengths. If x is the length of the path in feet, the number of colored tiles needed to make it is calculated using the rule (6x2y3)(4xy3) where y represents the number of white tiles. Which simplified expression represents the number of colored tiles used for a path of length x feet? 24x2y6 24x3y6 24x2y9 24x3y9

OpenStudy (anonymous):

i kno im on module 10 too im taking the test for the regular class so cant really help you

OpenStudy (anonymous):

ohh ok me 2

OpenStudy (radar):

What was your answer for 4.06?

OpenStudy (anonymous):

the last one

OpenStudy (radar):

O.K.